OpenStudy (anonymous):

the y intercept of the line tangent to y=x sin x at x=pi is... i don't understand this question :/

OpenStudy (danjs):

what course are you taking ?

OpenStudy (anonymous):

calculus

OpenStudy (danjs):

ok, the slope of the line tangent at that point of the graph, will he y ' (pi) First , find \[\frac{ dy }{ dx }=\frac{ d }{ dx }[x*\sin(x)]\]

OpenStudy (anonymous):

ok the derivative is x cos x sinx

OpenStudy (anonymous):

+ sinx i mean

OpenStudy (danjs):

The product rule, first times the derivative of the second, plus, the second times the derivative of the first. \[\frac{ dy }{ dx } = x *\frac{ d }{ dx }\sin(x) + \frac{ d }{ dx }[x] * \sin(x)\]

OpenStudy (danjs):

\[\frac{ dy }{ dx } = x*\cos(x) + (1) *\sin(x)\]

OpenStudy (anonymous):

yeah i understand that part :)

OpenStudy (danjs):

What is \[\frac{ dy }{ dx }\]when x=pi

OpenStudy (anonymous):

uhm -pi ?

OpenStudy (danjs):

The slope of the tangent line at x=pi will be \[\pi*\cos(\pi) + \sin(\pi)\]

OpenStudy (anonymous):

yes which is -pi right

OpenStudy (danjs):

right

OpenStudy (danjs):

the tangent line at x=pi will be \[y(x) = x*\sin(x)\] \[y - y(\pi) = y~'~(\pi)*(x-\pi)\]

OpenStudy (danjs):

y(pi) = 0 y - 0 = -pi*(x-pi)

OpenStudy (danjs):

\[y = -\pi*x + \pi^2\]

OpenStudy (danjs):

that is the equation for the line tangent to y(x) at x=pi

OpenStudy (anonymous):

ok so the y intercept is -pi?

OpenStudy (danjs):

The y-intercept can be found, by setting x=0 \[y=\pi^2\] ; x=0 the point

OpenStudy (danjs):

y-intercept \[(0,\pi^2)\]

OpenStudy (anonymous):

okay so how did u get pi^2?

OpenStudy (danjs):

one sec i have the graph...

OpenStudy (danjs):

using th point slope form for a line... y-y1 = m*(x-x1) the slope m is the derivative evaluated at x=pi ...m=-pi x1 = pi y1 = y(pi) = 0 the equation for the tangent line at x=pi is y = -pi*x + pi^2

OpenStudy (danjs):

When x=0, y=pi^2 that is the y intercept of the line

OpenStudy (danjs):

pi^2 is about 9.87

OpenStudy (anonymous):

wait i thought the y intercept was pi^2

OpenStudy (danjs):

it is, the graph i made has it in a decimal,, pi^2 = about 9387

OpenStudy (danjs):

9.87

OpenStudy (anonymous):

ohhh okay that makes sense

OpenStudy (danjs):

The tangent line to a graph y(x) at x=a , is found by \[y - y(a) = y~'(a)*(x-a)\]

OpenStudy (danjs):

just the point slope formula from algebra, with the slope being the derivative

OpenStudy (danjs):

point (a , y(a)) slope y '(a)

OpenStudy (anonymous):

okay thanks. I think i get it the only thing I'm a little confused about is exactly how to get the y intercept..Like what were the steps u took to get pi^2

OpenStudy (danjs):

If you are given a function f(x). To get the equation for a line tangent at some point x=a to the graph. you use the point slope formula from algebra. point (a , f(a)) slope f ' (a) ------------ point - slope formula for a line y - f(a) = f ' (a) *(x - a)

OpenStudy (danjs):

Once you have the equation for the line tangent at x=a. you let x=0, to get the y-intercept

OpenStudy (danjs):

(0 ,y) - y intercept

OpenStudy (danjs):

The equation for this line was: point (pi , y(pi)) = (pi , 0) slope = y '(pi) = - pi --------------------- Line point-slope form \[y - 0 = -\pi*(x-\pi)\]\[y = -\pi*x + \pi^2\]

OpenStudy (danjs):

When x = 0, y = pi^2 \[y = -\pi*0 +\pi^2\] the point \[(0,\pi^2)\]is the y-intercept of the tangent line to the graph at x=pi

OpenStudy (anonymous):

ohh okayy i get it

OpenStudy (danjs):

cool, i have time for another practice prob if you want to do another

OpenStudy (anonymous):

okay thanks. uhm another one i have is : for what values of x is f(x)=|x+1| differentiable? a) for all x can't equal 0, b) for all x is greater than or equal to 0 c) for all real numbers d)for all x can't equal 1 e)for all x can't equal -1

OpenStudy (danjs):

A graph is not differentiable where there is a sharp point.

OpenStudy (danjs):

That is the graph of the absolute value of x, shifted to the left 1 unit

OpenStudy (danjs):



OpenStudy (anonymous):

oh okay so x can't equal -1 ?

OpenStudy (danjs):

right, it is not differentiable when x=-1 there is no specific way to draw a tangent line at that point

OpenStudy (danjs):



OpenStudy (danjs):

lol, cant do it

OpenStudy (anonymous):

okay yeah i see what u mean. thank u so much ! i really appreciate it :)